OpenStudy (anonymous):

Simplifiy the quotient 4.5 x 10^3 / 9 x 10^7 Write the answer in scientific notation.

OpenStudy (jhannybean):

Is your question \[\frac{4.5\cdot 10^3}{9 \cdot 10^7}\]

OpenStudy (anonymous):

YEs

OpenStudy (jhannybean):

Alright, just expand your scientific notation

OpenStudy (anonymous):

idk what that means lol

OpenStudy (jhannybean):



OpenStudy (jhannybean):